The diagonal of a rectangle is 1m less than twice the length of one of the sides. The other side is 5m long. What is the diagonal
Answer by nerdybill(7384)

The diagonal of a rectangle is 1m less than twice the length of one of the sides. The other side is 5m long. What is the diagonal
.
Let d = diagonal
then
d/2 = length of one side
.
applying Pythagorean theorem:
d^2 = (d/2)^2 + 5^2
d^2 = d^2/4 + 25
4d^2 = d^2 + 100
3d^2 = 100
d^2 = 100/3
d = sqrt(100/3)
d = sqrt(100)/sqrt(3)
d = 10/sqrt(3)
d = 10/sqrt(3) * sqrt(3)/sqrt(3)
d = 10sqrt(3)/3 meters